OpenStudy (anonymous):

Use the formula for compound interest to find A A =p(1+r) n p= 30,000 r= 0.10 n (the expression) = 4 years I got $48,415. Can someone tell me if I am right or wrong and if so how i can fix my answer

OpenStudy (anonymous):

ur answer isn't correct please check ur calculation

OpenStudy (anonymous):

I believe you may have substituted the wrong numbers.

OpenStudy (anonymous):

reqd A=30000x(1+0.10)^4=30000*1.1*1.1*1.1*1.1 =30000*1.4641=43923

OpenStudy (anonymous):

30000(1+0.10)^4=$43923

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Okay I think I got confused when I figuring out the expression/scientific notation portion

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome